OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (anonymous):

A perpendicular line will have a slope that is the opposite reciprocal of the current slope. So the slope of the new line will be -1/4. Using the new slope and the point given, substitute into the point slope formula y=mx+b and solve for b -3=-1/4(2)+b

OpenStudy (anonymous):

b=7/2 so the equation is y=-1/4x+7/2 BUT--you said standard form, Ax+By=C y=-1/4x+7/2 1/4x+y=7/2 multiply everything by 4 to eliminate the fraction 4(1/4x+y=7/2) x+4y=14